OpenStudy (anonymous):

Match each equation with its y-intercep 5x-6y=30 A.-6 B.-5 C.5 D.6

OpenStudy (anonymous):

to find y intercept \[5(0)-6y=30\] \[-6y=30\] \[y=30/-6\] \[y=-5\]

OpenStudy (radha):

in eqn of form ax+by+c=0 y intercept= -constant term/coefficient of y= -c/b=-(-30)/-6=-5